OpenStudy (priyar):

How to calculate the displacement of a wave which undergoes damped oscillation due to an external periodic force?

OpenStudy (vincent-lyon.fr):

Please post full question.

OpenStudy (priyar):

i want to know how to derive the eq for a general case..

OpenStudy (priyar):

For ex: if there is a spring-block system and \[ F(t)= F _{0}\cos \omega _{d}t\] is the external periodic force and f=-bv is the damping force then how to find displacement as a function of time..?

OpenStudy (priyar):

@ParthKohli @samigupta8 @Vincent-Lyon.Fr

OpenStudy (priyar):

@inkyvoyd @ganeshie8

OpenStudy (priyar):

can u tell me the proper expression that we get at the end of this derivation?

OpenStudy (samigupta8):

D2x/dt^2

OpenStudy (priyar):

yes ofcourse!

OpenStudy (priyar):

atleast tell me what will be the end result..

Parth (parthkohli):

if you're given a bunch of options, just plug all of them and see which satisfies your equation.

OpenStudy (priyar):

what if i am given values..

OpenStudy (priyar):

\[m\frac{ d^2x }{ dt^2 } = -kx-bv+F _{0}\cos \omega _{d}t\]

OpenStudy (priyar):

then x=?

OpenStudy (vincent-lyon.fr):

x will be a sine function after a few oscillations , so plug in something like: \(x(t) = A\cos \omega _{d}t\)

OpenStudy (vincent-lyon.fr):

Then you will have to use a phasor diagram to add all the terms of the equation.

OpenStudy (irishboy123):

solve it as you would any 2nd order DE forst complementary solution for the homogeneous DE, ie where LHS = 0 then go after particular solution which you can get in a number of ways, but Vincent's suggestions is easiest and reflects practical reality, so guess a sinusoid form and plug it in

OpenStudy (vincent-lyon.fr):

I should have written : \(x(t) = A\cos (\omega _{d}t+\phi)\)

OpenStudy (priyar):

u told x will be a sine function but u have written cos funct. ..why? and what will be A in terms of \[A _{o}\] (highest amplitude/ the amplitude when there is no damping force..)

OpenStudy (irishboy123):

the particular solution to \(m\ddot x+ b\dot x+ kx = F _{0}\cos \omega _{d}t\) can be obtained in a number of ways, with probs the most basic being to assume a solution in the form \(y_p = \alpha \cos \omega _{d}t + \beta \sin \omega _{d}t\), with \(\alpha, \beta\) = const. so you stuff that into the actual DE and then get \(\alpha, \beta\) from balancing terms on LHS and RHS or you can combine the \( \alpha \cos \omega _{d}t + \beta \sin \omega _{d}t\) into a single cos or sin function with a phase shift. so \(x_p= A\cos (\omega _{d}t+\phi)\) or \(x_p= A\sin (\omega _{d}t+\theta)\) it's just this: http://www.sosmath.com/trig/addform/addform.html and can be seen here: http://hyperphysics.phy-astr.gsu.edu/hbase/oscdr.html

OpenStudy (priyar):

@IrishBoy123 can u pls tell me what is 'w' in that eq in the second link?

OpenStudy (priyar):

@IrishBoy123 i saw that link already..

OpenStudy (priyar):

i have a doubt in that... in the final expression what is 'w' ?

OpenStudy (priyar):

\[m(-Aw _{d}^2\sin(w _{d}t+\theta)) +b(Aw _{d}\cos(w _{d}t+\theta)) +kx = F _{0}cosw _{d}t\]

OpenStudy (priyar):

how to solve further? @Vincent-Lyon.Fr @IrishBoy123

OpenStudy (baru):

i think it will be simpler if you substitute \(x_p = \alpha \cos \omega _{d}t + \beta \sin \omega _{d}t\)... substitute it into the differential equation, and re-arrange it so you get something like \( ( ~~) \cos \omega _{d}t + (~~) \sin \omega _{d}t=F_0cos\omega_dt\) compare the coefficients of sin and cos terms of LHS and RHS to find alpha and beta

OpenStudy (vincent-lyon.fr):

Here is a way to do it if you are not familiar with complex numbers. Let's look for a solution as \(x(t)=A\cos(\omega t)\) responding to a driving force \(F(t)=F_o\cos(\omega t+\phi)\) We will need velocity and acceleration. \(v(t)=-\omega A\sin(\omega t)=\omega A\cos(\omega t+\pi /2)\) \(a(t)=-\omega ^2 A\cos(\omega t)=\omega ^2A\cos(\omega t+\pi )\) The differential equation can be drawn as a phasor:  Hence \(F_0=A\sqrt {(k-m\omega ^2)^2+(b\omega)^2}\) \(\tan \phi =\dfrac{b\omega}{k-m\omega ^2}\)

OpenStudy (vincent-lyon.fr):

Slight mistake corrected here: 

OpenStudy (priyar):

How did u draw that phasor diagram..? can u teach me how to draw one for the below equations..? \[F(t)=F _{0}\cos(\omega _{d}t)\]\[x(t)= Asin(\omega _{d}t +\phi)\]\[v(t)=A \omega _{d} cos(\omega _{d}t+\phi) \]\[a(t) = -A \omega _{d}^2\sin(\omega _{d}t+\phi) \]